OpenStudy (anonymous):

Find the derivative of y=x^(3+e), where e is a constant

OpenStudy (anonymous):

would it be (3+e)x^(2+e)?

OpenStudy (amistre64):

looks good; lets check that exponent tho to be safe. (3+e) -1 = 3+e-1 = 2+e ... looks right

OpenStudy (anonymous):

looks ok to me but don't forget about the.... + c ... on the end.

OpenStudy (anonymous):

+c?

OpenStudy (anonymous):

You don't need a +C because its a derivative not an integral.

OpenStudy (amistre64):

Rhino is confusing derivative for integration

OpenStudy (anonymous):

yeah my bad...forget that...sorry

OpenStudy (anonymous):

;)